OpenStudy (anonymous):

Solve the equation if each variable represents a whole number. 4x+y=15

OpenStudy (anonymous):

@Jerelb3

OpenStudy (anonymous):

4x+y=15 -4x -4x y=-4x+15

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

was it right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is it asking for what x and y is

OpenStudy (anonymous):

know it wants to know values of x and y that will make the open sentence true

OpenStudy (anonymous):

oh ok hold plz

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Well x=3 and y=3

OpenStudy (anonymous):

yeah so thats (3,3) thats one but they want me to find all the possible solutions.

OpenStudy (anonymous):

and how do i show my work like how do i show i got (3,3)?